# pEBR: A Probabilistic Approach to Embedding Based Retrieval

### v2

@@ Abstract
Embedding retrieval systems learn a shared semantic representation space for queries and items, enabling efficient retrieval through an approximate nearest-neighbor search. However, current industrial implementations face a critical limitation: using a fixed retrieval cutoff for all queries inevitably compromises performance, yielding insufficient recall for high-frequency (head) queries and reduced precision for low-frequency (tail) queries. This persistent challenge stems fundamentally from the frequentist paradigms dominating existing loss function designs. In this work, we introduce a novel framework, probabilistic Embedding-Based Retrieval (pEBR): maximum likelihood estimation-based and contrastive estimation-based, which learns the underlying probability distribution of relevant items for each query, compute adaptive cosine similarity cutoffs via probabilistic cumulative distribution functions (CDF), and automatically adapts to the distinct characteristics of head vs. tail queries. Experiments and ablation studies demonstrate that pEBR simultaneously improves precision and recall while maintaining computational efficiency.
@@
footnotetext: Corresponding author.
@@ 1 Introduction
Dense retrieval is the core technology in search (Liu et al., 2021; Zhao et al., 2024; Bao et al., 2024) and recommendationCovington et al. (2016); Li et al. (2020); Xing et al. (2025). The classic approaches are DSSM-like models (Huang et al., 2013; Zhang et al., 2020; Qiu et al., 2022; Li et al., 2021; Wang et al., 2023; Bao et al., 2024) etc., where query and item are embedded into dense vectors in the same semantic space. The relevant items are then retrieved by calculating the cosine similarity between the query vectors and the item vectors.
However, we notice that all the existing works on embedding-based retrieval do not pay enough attention to the model training algorithm, especially the loss function, which in practice causes some tricky problems as follows:
1) Ignores Queries Popularity Bias. Current models treat all queries equally, relying on simplistic heuristics, such as retrieving a fixed number of items or applying a fixed cosine similarity threshold. Despite well-documented performance differences between head, torso, and tail queries. Thus, incorporating query popularity could significantly improve retrieval quality. Specific Queries (e.g., "iPhone 16"): A small fixed cutoff misses relevant items, while a large one introduces noise. General Queries (e.g., "gift"): A fixed threshold cannot adapt to the wide variability in relevant items across categories (e.g., foods, clothes, cosmetics).
2) Lack of Probabilistic Foundation Limits Generalization. Without explicit assumptions about the item distribution given a query, models struggle to generalize, especially for long-tail queries where training data is sparse, degrading retrieval performance.
Before we introduce our new approach to address the above shortcomings, let’s first review the previous related works: probabilistic models, or more precisely probabilistic graphical models, which have a long history of successes in machine learning (Murphy, 2012; Koller and Friedman, 2009). There are a few models that significantly changed the landscape of machine learning in the past a few decades, such as Baum-Welch algorithm for training (McCallum, 2004), LDA (Blei et al., 2003), and probabilistic modeling with neural network Nguyen et al. (2017); Malinin (2019); Kohl et al. (2018).
Inspired by the previous successes of probabilistic spirits, here we propose a redesign of the embedding based retrieval to address the above challenges by using probabilistic modeling. Our contributions and the overview of the following sections can be summarized as follows.
@@
  • We then introduce two novel probabilistic methods: maximum likelihood estimation-based and contrastive estimation-based; the latter includes two instances, ExpNCE and BetaNCE.
  • Comprehensive comparative experiments, we demonstrate the superior effectiveness of our approach. Notice that this probabilistic modeling is also successfully applied in industrial-scale embedding-based retrieval systems.
@@ 2.1 Problem Definition and Notations
where $\mathbf{v}_{q}\in\mathbb{R}^{n}$ and $\mathbf{v}_{d}\in\mathbb{R}^{n}$ denotes the embedding of query and item, respectively. $s(.,.)$ computes the final score between the query and the item, such as the cosine similarity function, equivalently, inner product between normalized vectors, i.e., $s(\mathbf{v}_{q},\mathbf{v}_{d})=\mathbf{v}_{q}^{\top}\mathbf{v}_{d}$. The objective is to select the top $k$ items from a pool of candidate items for each given query $\mathbf{v}_{q}$.
@@ 2.2 Existing Frequentist Approaches
Most existing methods follow the frequentist approach to the loss function designs for embedding retrieval, including point-wise loss and pair-wise loss.
@@ 2.2.1 Point-Wise Loss
The point-wise based method converts the original retrieval task into a binary classification, whose goal is to optimize the embedding space where the similarity between the query and its relevant item is maximized, while the similarity between the query and the irrelevant items is minimized. It usually adopts the sigmoid cross-entropy to train the model:
$)\displaystyle L_{\textrm{pointwise}}(\mathcal{D})=-\sum_{i}\log\sigma(s(\mathbf{v}_{q_{i}},\mathbf{v}_{d^{+}_{i}}$))
$+\sum_{ij}\log\sigma(s(\mathbf{v}_{q_{i}},\mathbf{v}_{d_{ij}^{-}})),$
where $d^{+}_{i}$ denotes the relevant items for query $q_{i}$ and $d_{ij}^{-}$ denotes the irrelevant ones, $\sigma(x)=1/(1+\exp(-x))$ is the standard sigmoid function.
@@ 2.2.2 Pair-Wise Loss
This kind of method aims to learn the partial order relationship between positive and negative items from the perspective of metric learning. The classical work contains triple loss Li et al. (2023), margin lossLiu et al. (2017), A-Softmax loss, and several variants Wang et al. (2018). Without loss of generality, we formulate the loss as softmax cross-entropy: $L_{\textrm{pairwise}}(\mathcal{D})$ $=-\sum_{i}\log\left(\frac{\exp(s(\mathbf{v}_{q_{i}},\mathbf{v}_{d^{+}_{i}})/\tau)}{\sum_{j}\exp(s(\mathbf{v}_{q_{i}},\mathbf{v}_{d_{ij}})/\tau)}\right),$ where $\tau$ is the so-called temperature parameter: a lower temperature is less forgiving of mis-prediction of positive items by the model. The larger the value, the more concerned it is with the negative item.
@@ 3 Method
We explore the probabilistic approach, which contains the choice of the item distribution function and our carefully designed loss function based on the item distribution.
@@ 3.2.1 Parametric Exponential InfoNCE (ExpNCE)
A nice property of the above probabilistic modeling is that we can use a simple cumulative density function (CDF) to decide the cutoff threshold. The CDF for the above truncated exponential distribution can be derived as follows. $\mathbb{P}_{\rm{ExpNCE}}(x<t)=\frac{\int_{-1}^{t}e^{x/\tau_{q}}dx}{\int_{-1}^{1}e^{x/\tau_{q}}dx}=\frac{e^{t/\tau_{q}}-e^{-1/\tau_{q}}}{e^{1/\tau_{q}}-e^{-1/\tau_{q}}},$ which can be easily computed.
@@ 3.2.2 Parametric Beta InfoNCE (BetaNCE)
We call a probability distribution compactly supported if its cumulative distribution function $F$ satisfies $F((-\infty,-x])=1-F([x,\infty))=0$ for some $x>0$. In other words, all the mass is contained in the finite interval [x,x][-x,x]. The best-known family of compact distributions in statistics is probably the Beta distributions, whose pdf are defined on $[0,1]$. Since the cosine similarity used in two-tower model has value range of [1,1][-1,1], we expand the definition range of Beta distributions to [1,1][-1,1] and define its density as
@@
One difficulty working with Beta distributions is that its CDF has no closed form, but rather is given by the incomplete Beta function [^1] https://www.tensorflow.org/api_docs/python/tf/math/betainc: $B(t;\alpha,\beta)=\int_{0}^{t}x^{\alpha-1}(1-x)^{\beta-1}dx$ is the incomplete Beta integral. The CDF for the above Beta distribution can be derived as follows. $\mathbb{P}_{\rm{BetaNCE}}(x<t)$ $=\frac{\int_{-1}^{t}((1+x)/2)^{\alpha-1}((1-x)/2)^{\beta-1}dx}{\int_{-1}^{1}((1+x)/2)^{\alpha-1}((1-x)/2)^{\beta-1}dx}$ $=\frac{\int_{0}^{\frac{t+1}{2}}x^{\alpha-1}(1-x)^{\beta-1}dx}{\int_{0}^{1}x^{\alpha-1}(1-x)^{\beta-1}dx}$ $=\frac{\Gamma(\alpha+\beta)}{\Gamma(\alpha)\Gamma(\beta)}B(\frac{t+1}{2};\alpha,\beta).$
@@ 4 Experiments
In this section, we first explain the details of the dataset, experimental setup, baseline methods and evaluation metrics. Then we show the comparison of experimental results between the baseline methods and pEBR. Finally we show the ablation study to intuitively illustrate how pEBR could perform better.
@@ 4.1.1 Dataset and setup
Our model was trained on a randomly sampled dataset consisting of 87 million user click logs collected over a period of 60 days. We trained the model on an Nvidia GPU A100 and employed the AdaGrad optimizer with a learning rate of 0.05, batch size of 1024, and an embedding dimension of 128. The training process converged after approximately 200,000 steps, which took around 12 hours.
@@ 4.1.2 Comparative Backbone
  • DSSM-topk refers to the method that cuts off retrieved items using a fixed threshold of numbers.
  • DSSM-score cuts off retrieved items with a fixed threshold of relevance score. Items with relevance scores below the threshold are discarded.
  • pEBR cuts off retrieved items with a threshold derived by a probabilistic model, specifically the CDF value of the learned item distribution. We are using a default CDF value $0.5$ if not further specified. In this experiment, we focus on using BetaNCE with $\beta=1$ as shown in Equation (3) to demonstrate the effectiveness of the methods. But our method is general enough to accommodate other distributions, such as ExpNCE.
As our method is model-agnostic and can be applied to any two-tower architecture, we also conducted experiments with a BGE Xiao et al. (2024) (transformer-based) backbone to verify this compatibility. Details of these additional experiments are presented in Appendix B.
@@ 4.1.3 Experimental metrics
We use two widely reported metrics, Recall@$k$ and Precision@$k$, to evaluate the quality and effectiveness of retrieval models (Zhang et al., 2021; Zhu et al., 2018; Yuan et al., 2023). There are some nuances that need a little more explanation. In DSSM-topk, a fixed number $k$ of items is retrieved for each query to calculate R@$k$ and P@$k$. While in DSSM-score and pEBR, a threshold of relevance score or a CDF value is used to cut off retrieved items, which results in a different number of items for each query. Thus for fair comparison, we tune the threshold slightly to make an average number $k$ of items to be retrieved for all queries to calculate the metrics.
Moreover, there is a tradeoff between precision and recall when varying the $k$ value. In the experiments, we choose $k=1500$ to optimize for recall, since it is the main goal of a retrieval system. Thus, the precision value is relatively very low.
@@ 4.2 Experimental Results
Since the retrieval performance differs significantly from head, torso to tail queries, we implement tripartite dataset partitioning for stratified evaluation.
As shown in Table 1 we can draw the following conclusions:
@@ 4.3 Online A/B Testing
Through a continuous and effective week-long controlled experiment, from Table 2 we can discover that ours showed improvements over the previous base model, with a 0.19% boost in User Conversion Rate (UCVR) (p-value < 0.02), translating into substantial revenue gains for the company.
Figure 2: Histogram of the number of retrieved items for head, torso, and tail queries, where we fix the CDF value to be 0.985.
@@ 4.4.2 Dynamic Retrieval Effect
To investigate the effect of dynamic retrieval, we analyze the retrieval results from two perspectives. On one hand, we examine the distribution of the number of retrieved items under a fixed probabilistic CDF threshold. On another hand, we evaluate the number of retrieved items under different probabilistic CDF thresholds. These analyses allow us to gain insight into how dynamic retrieval influences the retrieval results.
In Figure 2, we show the histogram of retrieved item numbers cut off by CDF value 0.985 for both head, torso, and tail queries. In detail, we filter out the items with the cosine similarity $x$ bellowing a threshold $t$ to ensure the equation $\mathbb{P}(x>=t)=0.985$. In general, head queries have a more uniform distribution and the numbers lie mostly in the range of [500, 1500], while torso and tail queries share a similar skewed distribution and the numbers lie in the range of [0, 1000] and [0, 500], respectively. This indicates that head queries retrieve more items than torso and tail queries, which confirms the assumption that queries with higher popularity need more candidates and queries with lower popularity need fewer.
In Table 3, we present the average number of retrieved items under different CDF values with the same cutting off strategy in Figure 2. As the probability decreases, queries claim a higher relevance level for items, thus the average numbers of retrieved items decreases accordingly. Comparing different queries, head queries retrieve more items than torso and tail queries under all CDF thresholds, which again confirms the conclusion above.

### v3

@@ Abstract
Embedding-based retrieval aims to learn a shared semantic representation space for both queries and items, enabling efficient and effective item retrieval through approximate nearest neighbor (ANN) algorithms. In current industrial practice, retrieval systems typically retrieve a fixed number of items for each query. However, this fixed-size retrieval often results in insufficient recall for head queries and low precision for tail queries. This limitation largely stems from the dominance of frequentist approaches in loss function design, which fail to address this challenge in industry. In this paper, we propose a novel probabilistic Embedding-Based Retrieval (pEBR) framework. Our method models the item distribution conditioned on each query, enabling the use of a dynamic cosine similarity threshold derived from the cumulative distribution function (CDF) of the probabilistic model. Experimental results demonstrate that pEBR significantly improves both retrieval precision and recall. Furthermore, ablation studies reveal that the probabilistic formulation effectively captures the inherent differences between head-to-tail queries.
@@
footnotetext: Corresponding authors.
@@ 1 Introduction
Embedding retrieval is the core technology in search (Liu et al., 2021; Zhao et al., 2024; Bao et al., 2024) and recommendationCovington et al. (2016); Li et al. (2020); Xing et al. (2025). The classic approaches are DSSM-like models (Huang et al., 2013; Zhang et al., 2020; Qiu et al., 2022; Li et al., 2021; Wang et al., 2023; Bao et al., 2024) etc., where the query and item are embedded into dense vectors in the same semantic space. The relevant items are then retrieved by calculating the cosine similarity between the query vectors and the item vectors.
However, we notice that existing works on embedding-based retrieval do not pay enough attention to the model training algorithm, especially the loss function, which in practice causes some tricky problems as follows: 1) Heuristic and suboptimal retrieval cutoffs. Existing embedding-based retrieval systems rely on heuristic retrieval cutoffs, such as a fixed number of items or a fixed cosine similarity threshold. This leads to imbalanced retrieval performance: head queries suffer from low recall, while tail queries suffer from low precision. For instance, "iPhone 16" is a very specific query with few relevant items, while "gift" is broad, covering many products across categories. Thus, one can easily find that retrieving a fixed number of items would be suboptimal here, since if the number is small we will miss many relevant products for "gift" and if the number is large we will introduce many irrelevant items for "iPhone 16". On the other hand, it is also suboptimal to use a fixed cosine threshold, since all the existing model training algorithms have never taken that into account in their design. 2) Ignoring Query Popularity Bias. Current two-tower training algorithms treat all queries equally and lack a probabilistic formulation to model item relevance distributions. As a result, they fail to account for query popularity differences (head vs. tail queries) and struggle to generalize, especially for long-tail queries with limited training data.
Before we introduce our new approach to address the above shortcomings, let us first review previous related works: probabilistic models, or more precisely probabilistic graphical models, which have a long history of success in machine learning (Murphy, 2012; Koller and Friedman, 2009). Several models have significantly changed the landscape of machine learning in recent decades, such as the Baum-Welch algorithm for training (McCallum, 2004), LDA (Blei et al., 2003), and probabilistic modeling with neural network Nguyen et al. (2017); Malinin (2019); Kohl et al. (2018).
Inspired by the previous successes of probabilistic spirits, here we propose a redesign of the embedding-based retrieval to address the above challenges by using probabilistic modeling. Our contributions and the overview of the following sections can be summarized as follows.
@@
  • We then propose two novel probabilistic methods: one based on maximum likelihood estimation and another based on contrastive estimation; the latter includes two instances, ExpNCE and BetaNCE.
  • We conduct both offline and online experiments to demonstrate the effectiveness of our model, and perform ablation studies to help understand how the model works.
@@ 2.1 Problem Definition and Notations
where $\mathbf{v}_{q}\in\mathbb{R}^{n}$ and $\mathbf{v}_{d}\in\mathbb{R}^{n}$ denotes the embedding of the query and the item, respectively. $s(.,.)$ computes the final score between the query and the item, such as the cosine similarity function, equivalently inner product between normalized vectors, i.e., $s(\mathbf{v}_{q},\mathbf{v}_{d})=\mathbf{v}_{q}^{\top}\mathbf{v}_{d}$, where the superscript denotes matrix transpose. The objective is to select the top $k$ items from a pool of candidate items for each given query $\mathbf{v}_{q}$.
@@ 2.2 Existing Frequentist Approaches
Most existing embedding retrieval adopt a frequentist approach to loss function design, which can be categorized into two types: point-wise loss and pair-wise loss. We discuss each category in detail.
@@ 2.2.1 Point-Wise Loss
The point-wise based method converts the original retrieval task into a binary classification, whose goal is to optimize the embedding space where the similarity between the query and its relevant item is maximized, while the similarity between the query and the irrelevant items is minimized. It usually adopts the sigmoid cross-entropy to train the model: $)\displaystyle L_{\textrm{pointwise}}(\mathcal{D})=-\sum_{i}\log\sigma(s(\mathbf{v}_{q_{i}},\mathbf{v}_{d^{+}_{i}}$)) $+\sum_{ij}\log\sigma(s(\mathbf{v}_{q_{i}},\mathbf{v}_{d_{ij}^{-}})),$ where $d^{+}_{i}$ denotes the relevant items for query $q_{i}$ and $d_{ij}^{-}$ denotes the irrelevant ones, $\sigma(x)=1/(1+\exp(-x))$ is the standard sigmoid function.
@@ 2.2.2 Pair-Wise Loss
This kind of method aims to learn the partial order relationship between positive and negative items from the perspective of metric learning. The classical work contains triple loss, margin loss, A-Softmax loss, and several variants (margin angle cosine). Without loss of generality, we formulate the loss as softmax cross-entropy:
$L_{\textrm{pairwise}}(\mathcal{D})$ $=-\sum_{i}\log\left(\frac{\exp(s(\mathbf{v}_{q_{i}},\mathbf{v}_{d^{+}_{i}})/\tau)}{\sum_{j}\exp(s(\mathbf{v}_{q_{i}},\mathbf{v}_{d_{ij}})/\tau)}\right),$
where $\tau$ is the so-called temperature parameter: a lower temperature is less forgiving of mis-prediction of positive items by the model. In the same direction, researchers later proposed more advanced loss functions, by introducing max margin in angle space (Liu et al., 2017), in cosine space (Wang et al., 2018) and so on.
@@ 3 Method
In this section, we explore the probabilistic approach, which contains the choice of the item distribution function and our carefully designed loss function based on the item distribution.
@@ 3.2.1 Parametric Exponential InfoNCE (ExpNCE)
A nice property of the above probabilistic modeling is that we can use a simple cumulative density function (CDF) to decide the cutoff threshold. The CDF for the above truncated exponential distribution can be derived as follows.
$\mathbb{P}_{\rm{ExpNCE}}(x<t)$ $=\frac{\int_{-1}^{t}e^{x/\tau_{q}}dx}{\int_{-1}^{1}e^{x/\tau_{q}}dx}$
$=\frac{e^{t/\tau_{q}}-e^{-1/\tau_{q}}}{e^{1/\tau_{q}}-e^{-1/\tau_{q}}},$
which can be easily computed.
@@ 3.2.2 Parametric Beta InfoNCE (BetaNCE)
We call a probability distribution compactly supported if its cumulative distribution function $F$ satisfies $F((-\infty,-x])=1-F([x,\infty))=0$ for some $x>0$. In other words, all the mass is contained in the finite interval [x,x][-x,x]. The best-known family of compact distributions in statistics is probably the Beta distributions, whose pdf are defined on $[0,1]$. Since the cosine similarity used in the two-tower model has a value range of [1,1][-1,1], we expand the definition range of Beta distributions to [1,1][-1,1] and define its density as
@@
One difficulty working with Beta distributions is that its CDF has no closed form, but rather is given by the incomplete Beta function [^1] https://www.tensorflow.org/api_docs/python/tf/math/betainc: $B(t;\alpha,\beta)=\int_{0}^{t}x^{\alpha-1}(1-x)^{\beta-1}dx$ is the incomplete Beta integral. The CDF for the above Beta distribution can be derived as follows.
$\mathbb{P}_{\rm{BetaNCE}}(x<t)$ $=\frac{\int_{-1}^{t}((1+x)/2)^{\alpha-1}((1-x)/2)^{\beta-1}dx}{\int_{-1}^{1}((1+x)/2)^{\alpha-1}((1-x)/2)^{\beta-1}dx}$
$=\frac{\int_{0}^{\frac{t+1}{2}}x^{\alpha-1}(1-x)^{\beta-1}dx}{\int_{0}^{1}x^{\alpha-1}(1-x)^{\beta-1}dx}$
$=\frac{\Gamma(\alpha+\beta)}{\Gamma(\alpha)\Gamma(\beta)}B(\frac{t+1}{2};\alpha,\beta).$
@@ 4 Experiments
In this section, we first explain the details of the dataset, experimental setup, baseline methods and evaluation metrics. Then we show the comparison of experimental results between the baseline methods and our proposed method pEBR. Finally we show the ablation study to intuitively illustrate how pEBR could perform better.
@@ 4.1.1 Dataset and Setup
Our model was trained on a randomly sampled dataset consisting of 87 million user click logs collected over a period of 60 days. We trained the model on an Nvidia GPU A100 and employed the AdaGrad optimizer with a learning rate of 0.05, batch size of 1024, and an embedding size of 128. The training process converged after approximately 200,000 steps, which took around 12 hours.
@@ 4.1.2 Comparative Backbone
  • DSSM-topk refers to the method that cuts off items using a fixed threshold of number.
  • DSSM-score refers to the method that cuts off items with a fixed threshold of relevance score. Items with relevance scores below the threshold are discarded.
  • pEBR refers to our proposed probabilistic method which cuts off items with a threshold derived by a probabilistic model, specifically the CDF value of the learned item distribution. We are using a default CDF value $0.5$ if not further specified. In this experiment, we focus on using BetaNCE with $\beta=1$ as shown in Equation (3) to demonstrate the effectiveness of the methods. But our method is general enough to accommodate other distributions, such as ExpNCE.
As our method is model-agnostic and can be applied to any two-tower architecture, we also conduct experiments with a BGE Xiao et al. (2024) (transformer-based) backbone to verify this compatibility. Details of these additional experiments are presented in Appendix B.
@@ 4.1.3 Experimental Metrics
We use two widely reported metrics, Recall@$k$ and Precision@$k$, to evaluate the quality and effectiveness of retrieval models (Zhang et al., 2021; Zhu et al., 2018; Yuan et al., 2023). There are some nuances that need a little more explanation. In DSSM-topk, a fixed number $k$ of items is retrieved for each query to calculate R@$k$ and P@$k$. While in DSSM-score and pEBR, a threshold of relevance score or CDF value is used to cut off retrieved items, which results in a variable number of items for each query. To ensure a fair comparison, we slightly adjust the threshold so that the average number of retrieved items across all queries equals $k$ when computing the evaluation metrics.
Moreover, there is a tradeoff between precision and recall when varying the $k$ value. In the experiments, we choose $k=1500$ to optimize for recall, since it is the main goal of a retrieval system. Thus, the precision value is relatively low.
@@ 4.2 Experimental Results
Since retrieval performance varies considerably across head, torso, and tail queries, we adopt a tripartite dataset partitioning strategy for stratified evaluation. As shown in Table 1, we can draw the following conclusions:
@@ 4.3 Online A/B Testing
Through a continuous week-long controlled experiment, table 2 shows that our proposed method achieved consistant improvements over the previous baselines, yielding a 0.19% increase in User Conversion Rate (UCVR) (p-value < 0.02), translating into substantial revenue gains for the company.
Figure 2: Histogram of the retrieved item numbers for head, torso, and tail queries, with the CDF value set to 0.985.
@@ 4.4.2 Dynamic Retrieval Effect
To investigate the effect of dynamic retrieval, we analyze the retrieval results from two perspectives. On the one hand, we examine the distribution of the number of retrieved items under a fixed probabilistic CDF threshold. On the other hand, we evaluated the number of items recovered under different probabilistic CDF thresholds. These analyses allow us to gain insight into how dynamic retrieval influences the retrieval results.
In Figure 2, we show the histogram of retrieved item numbers cut off by the CDF value 0.985 for both head, torso, and tail queries. In detail, we filter out the items with the cosine similarity x bellowing a threshold t to ensure the equation P(x >= t) = 0.985. In general, head queries have a more uniform distribution, and the numbers lie mostly in the range of [500, 1500], while the torso and tail queries share a similar skewed distribution and the numbers lie in the range of [0, 1000] and [0, 500], respectively. This indicates that head queries retrieve more items than torso and tail queries, which confirms the assumption that queries with higher popularity need more candidates and queries with lower popularity need fewer.
In Table 3, we present the average number of retrieved items under different CDF values with the same cutting off strategy in Figure 2. As the probability decreases, queries claim a higher relevance level for items, and thus the average number of retrieved items decreases accordingly. Comparing different queries, head queries retrieve more items than torso and tail queries under all CDF thresholds, which again confirms the conclusion above.